OpenStudy (anonymous):

MEDALS ARE GIVEN!!!!!WHOEVER ANSWERS FIRST! a piece of tile is 4 1/2 inches long and 3 1/4 inches wide. What is the area of the tile?

OpenStudy (anonymous):

First convert to decimals 4 1/2 = 4.5 3 1/4 = 3.25 Then multiply the numbers together 4.5 * 3.25 = Answer

OpenStudy (anonymous):

ok thanks